### v2

# MA4DIV: Multi-Agent Reinforcement Learning for Search Result Diversification

Abstract.
The objective of search result diversification (SRD) is to ensure that selected documents cover as many different subtopics as possible. Existing methods primarily utilize a paradigm of ”greedy selection”, i.e., selecting one document with the highest diversity score at a time. These approaches tend to be inefficient and are easily trapped in a suboptimal state. In addition, some other methods aim to approximately optimize the diversity metric, such as $\alpha$-NDCG, but the results still remain suboptimal. To address these challenges, we introduce Multi-Agent reinforcement learning (MARL) for search result DIVersity, which called MA4DIV. In this approach, each document is an agent and the search result diversification is modeled as a cooperative task among multiple agents. This approach allows for directly optimizing the diversity metrics, such as $\alpha$-NDCG, while achieving high training efficiency. We conducted preliminary experiments on public TREC datasets to demonstrate the effectiveness and potential of MA4DIV. Considering the limited number of queries in public TREC datasets, we construct a large-scale dataset from industry sources and show that MA4DIV achieves substantial improvements in both effectiveness and efficiency than existing baselines on a industrial scale dataset.
Search Result Diversification; Learning to Rank; Multi-Agent Cooperation; Reinforcement Learning
1. Introduction
How to provide diverse search results to users is an important yet complicated problem in information retrieval. The search result diversification (SRD) (Agrawal et al., 2009) means that for a user’s search query, we need to provide search results that cover a wide range of subtopics to meet the requirements of users who may have multiple interpretations or intentions. Many diversity-aware metrics, such as $\alpha$-NDCG (Clarke et al., 2008), ERR-IA (Chapelle et al., 2011), S-recall (Zhai et al., 2015), have been proposed to evaluate the effectiveness of search result diversification. Since the diversity of a given document is affected by the documents that precede it, optimizing some of the diversity metrics is an NP-Hard problem (Carterette, 2011). Therefore, the diversity of search results has become an important and challenging research topic.
The typical search result diversification method includes heuristic methods and learning-based methods. These methods use ”greedy selection” to sequentially select documents to build a diversified ranking list in a step-by-step manner. At each step, the ranking model selects the best document for the current ranking position according to the additional utility it can bring to the whole ranking list. Different methods use different criteria in the greedy selection. Maximal marginal relevance (MMR) (Carbonell and Goldstein, 1998) algorithm proposes maximal marginal, which is defined as the sum of query-document dependencies, and takes the maximum document distance as the utility. xQuAD (Santos et al., 2010) is another widely used diverse ranking model which defines the utility to explicitly account for relationship between documents retrieved for the original query and the possible sub-queries. In recent years, many methods based on machine learning (Zhu et al., 2014; Xu et al., 2017; Li et al., 2009; Mihalkova and Mooney, 2009; Qin et al., 2023) have been proposed to solve the search result diversification task. For example, the relational learning to rank (R-LTR) (Zhu et al., 2014) and its variants utilize a linear combination of correlation features and novelty features.
All above methods focus on estimating the utility of candidate documents directly based on user queries and previously selected documents, but fail to explicitly model and fully utilized the utility that the user has gain in examining previous documents. Xia et al. (2017) propose the MDP-DIV framework to model the search result diversification ranking as Markov Decision Process (MDP) (Puterman, 2014), and uses a reinforcement learning algorithm (Williams, 1992) to optimize the ranking model. While MDP-DIV still use a ”greedy selection” to select next candidate result, it can easily lead to sub-optimal ranking list. $\mathrm{M^{2}Div}$ (Feng et al., 2018) introduces Monte Carlo Tree Search (MCTS) (Coulom, 2006) for high-quality exploration. This method helps to alleviate the problem of being stuck in a sub-optimal solution at a cost of increasing training and inference time.
Besides these greedy selection methods that sequentially score and select each candidate results, some recent studies DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) propose to simultaneously score all candidate documents and obtain the whole ranking list by sorting these scores in a descending order. This paradigm can greatly improve the inference efficiency as the documents can be scored in parallel. In addition, DALETOR and MO4SRD propose a differentiable proxy for the non-differentiable diversity metric $\alpha$-NDCG (Clarke et al., 2008), which enables a gradient-based optimization for the diversity metric. However, since the optimization goal is an approximation to the real evaluation metric, the solution may still be sub-optimal.
To summarize, although the problem of search result diversification has been extensively investigated, existing methods still have some limitations in terms of the effectiveness and efficiency. From the perspective of effectiveness, the heuristic-based methods (e.g. (Carbonell and Goldstein, 1998; Santos et al., 2010)) and the approximation-based methods (e.g.(Yan et al., 2021; Yu, 2022)) may be sub-optimal in optimizing the end diversity metric, such as $\alpha$-NDCG and ERR-IA. From the perspective of efficiency, those RL-based approaches face a intrinsic challenge of exploring a huge space of all possible rankings, which inevitably leads to sub-optimal solution (Xia et al., 2017) or high time-complexity in both training and inference (Feng et al., 2018).
In this paper, in order to alleviate the problems of existing methods mentioned above, we propose to formalize the diverse ranking as a multi-agent cooperation process. Since Multi-Agent reinforcement learning (MARL) algorithm is used to optimize the DIVersity ranking model, we call this new method MA4DIV. Specifically, by treating each document as an independent agent within a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) setting, we simulate a fully cooperative multi-agent task. In this setup, each agent (document) makes action selections based on observations that include the features of both the query and the documents, aiming to maximize a shared reward function that is directly related to the diversity evaluation metrics of search results. Considering permutation equivariant (Pang et al., 2020) property of the ranking task, our MA4DIV framework leverages the Multi-Head Self-Attention (MHSA) for capturing the intricate relationships between documents. Furthermore, MA4DIV employs value decomposition (Rashid et al., 2018) to optimize global diversity directly by the structures of mixing network and hypernetworks during training.
The advantages of our MA4DIV over existing works include:
(1) Compared to ”greedy selection” paradigm, MA4DIV simultaneously predicts the ranking scores of all documents, which can improve the efficiency of ranking process and alleviate the local optimum problem.
(2) The ranking scores of all documents and a ranked documents list with diversity can be obtained in one time step. Therefore, MA4DIV has higher exploration efficiency than single-agent RL algorithms, such as MDP-DIV and $\mathrm{M^{2}Div}$, which must go through an entire episode to explore a diversity ranking list.
(3) The training of the MA4DIV model does not require any approximations of different diversity metrics as they can be directly optimized as reward in the multi-agent reinforcement learning process.
In order to demonstrate the effectiveness and efficiency of MA4DIV, we conducted experiments on TREC benchmark datasets and Baidu dataset. The experimental results on TREC datasets showed that, MA4DIV achieved the state-of-the-art performance on some evaluation metrics, and training time was significantly shorter than baselines. Considering the small number of queries in TREC dataset (only 198 valid queries), we built a larger diversity dataset based on the real Baidu search engine data, which is called DU-DIV. On the DU-DIV dataset, MA4DIV achieved the state-of-the-art performance on all evaluation metrics, and demonstrated a substantial improvement in exploration efficiency and training efficiency.
2. RELATED WORK
2.1. Search Result Diversification
The construction of a diverse ranking list in search is commonly formalized as a sequential document selection process or a process of greedy sequential decision-making. The aim is to design effective criteria for estimating the utility a document can bring, which is believed to depend not only on the document itself but also on the preceding documents. Various models have been proposed to guide this selection.
Carbonell and Goldstein introduced the maximal marginal relevance criterion (Carbonell and Goldstein, 1998), which uses a linear combination of query-document relevance and document novelty to determine the document to be selected. This process is iterative, with the marginal relevance score updated as the number of documents in the result set increases. An extension of this concept, the probabilistic latent MMR model (Guo and Sanner, 2010), was proposed by Guo and Sanner.
Approaches such as xQuAD (Santos et al., 2010) directly model different aspects of a query, estimating utility based on the relevance of retrieved documents to identified sub-queries or aspects. Hu et al. (Hu et al., 2015) proposed a utility function that leverages hierarchical intents of queries, aiming to maximize diversity in the hierarchical structure. Other researchers, like He et al. (He et al., 2012), have proposed combining implicit and explicit topic representations to create better diverse rankings, while Gollapudi and Sharma (Gollapudi and Sharma, 2009) proposed an axiomatic approach to result diversification.
Machine learning techniques have also been applied to construct diverse ranking models, often adhering to the sequential document selection or greedy sequential decision-making framework. These techniques aim to automatically learn the utility function based on training queries. Several researchers have defined utility as a linear combination of handcrafted relevance and novelty features (Zhu et al., 2014; Xia et al., 2015; Xu et al., 2017). In (Xia et al., 2016), novelty can be modeled using deep learning models such as neural tensor networks. Radlinski et al. (2008) propose to learn a diverse ranking of documents directly based on users’ clicking behavior.
Reinforcement learning algorithms also are proposed to improve the effectiveness of ”greedy selection” paradigm in search result diversification. MDP-DIV (Xia et al., 2017) consider the diverse ranking process as a Markov Decision Process (MDP). In MDP-DIV, agent select one document in one time step, so a ranked list can be obtained after an episode. And the episode reward can be any metrics, such as $\alpha$-DCG. While Feng et al. (2018) think that ”greedy selection” paradigm is easily lead to local optimum, so $\mathrm{M^{2}Div}$ is proposed to introduce Monte Carlo Tree Search (MCTS) into MDP ranking process. In this way, $\mathrm{M^{2}Div}$ does alleviate the problem that tend to fall into local optimum by conducting high-quality exploration, but also brings expensive training and inference costs.
There are other methods which approximate the diversity evaluation metrics to be differentiable and optimize the approximated metrics directly through the loss function. Based on score-and-sort paradigm, DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) are proposed to derive differentiable diversification-aware losses which are approximation of different diversity metrics, such as $\alpha$-NDCG.
2.2. Reinforcement Learning for IR
Xu et al. (Wei et al., 2017) formulate a ranking process as an MDP and train the ranking model with a policy gradient algorithm of REINFORCE (Williams, 1992). Singh et al. (Singh and Joachims, 2019) propose PG Rank algorithm to enhance the fairness of ranking process with reinforcement learning methods. PPG (Xu et al., 2020) uses pairwise comparisons of two sampled document list with in a same query, which makes an unbiased and low variance policy gradient estimations. Yao et al. (Yao et al., 2020) model the interactions between search engine and users with a hierarchical Markov Decision Process, which improves the personalized search performance. RLIRank (Zhou and Agichtein, 2020) implement a learning to rank (LTR) model for each iteration of the dynamic search process. Zou et al. (2019) formulate the ranking process as a multi-agent Markov Decision Process, where the interactions among documents are considered in the ranking process. However, Zou et al. (2019) conduct the single-agent reinforcement learning algorithm REINFORCE to optimize the ranking model.
2.3. Multi-Agent Reinforcement Learning
Value decomposition (Sunehag et al., 2017; Rashid et al., 2018; Wang et al., 2020b, a; Chen et al., 2022b, a) and Actor-Critic (Lowe et al., 2017; Zhang et al., 2022, 2023a, 2023b; Yu et al., 2022) are two typical branches of multi-agent reinforcement learning (MARL). Among these, QMIX (Rashid et al., 2018) is the algorithm that first obtains the global utility function by nonlinear combination of individual utility functions, which can assign the global reward to each agent implicitly and nonlinearly. In this paper, we conduct the framework of QMIX as a reference to optimize the diverse ranking process.
3. Background
3.1. Dec-POMDP
In this work, we consider the process of diversity ranking as a fully cooperative multi-agent task which is usually modeled as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato, 2016). Dec-POMDP is formally defined as a tuple $G=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{Z},\mathcal{O},n,\gamma\right\rangle$. $s\in\mathcal{S}$ is the global state of the environment. Each agent $i\in\mathcal{G}\equiv\{1,\dots,n\}$ chooses an action $a^{i}\in A$ which forms the joint action $\textbf{a}\in\mathcal{A}\equiv A^{n}$. The state transition function is modeled as $\mathcal{P}(s^{\prime}|s,\textbf{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$. The reward function which is modeled as $r(s,\textbf{a}):\mathcal{S}\times\mathcal{A}$ is shared by all agents and the discount factor is $\gamma\in[0,1)$. It follows partially observable settings, where agents do not have access to the global state. Instead, it samples observations $z\in\mathcal{Z}$ according to observation function $\mathcal{O}(s,i):\mathcal{S}\times\mathcal{A}\to\mathcal{Z}$. Each agent has an action-observation history trajectory $\tau^{i}\in T\equiv(\mathcal{Z}\times\mathcal{A})^{*}$, on which it conditions a stochastic policy $\pi^{i}(u^{i}|\tau^{i}):T\times\mathcal{A}\to[0,1]$. In our algorithm, the joint policy $\pi$ is based on action-value function $Q^{\pi}_{tot}(s_{t},\textbf{a}_{t})=\mathbb{E}_{s_{t+1}:\infty,\textbf{a}_{t+1}:\infty}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}|s_{t},\textbf{a}_{t}]$. The final goal is to get the optimal action-value function $Q^{*}$.
3.2. Diversity Evaluation Metrics
In this section, we introduce three diversity evaluation metrics, namely $\alpha$-NDCG (Clarke et al., 2008), ERR-IA (Chapelle et al., 2011) and S-recall (Zhai et al., 2015). The basic setting is that there are $n$ documents in a query, and each document may cover 1 to $m$ subtopics.
3.2.1. $\alpha\textrm{-}NDCG$
Firstly, the $\alpha$ discounted cumulative gain ($\alpha\textrm{-}DCG$) is defined as Equation (1).
(1) $\alpha\textrm{-}DCG=\sum^{n}_{i=1}\sum^{m}_{l=1}\frac{y_{il}(1-\alpha)^{c_{li}}}{log_{2}(1+r_{i})}$
$y_{il}=1$ means that subtopic $l$ is covered by document $i$ and $y_{il}=0$ is the opposite. $\alpha$ is a parameter between 0 to 1, which quantifies the probability of a reader getting the information about a specific subtopic from a relevant document. $r_{i}$ is the ranking position of document $i$, and $c_{li}$ is the number of times that the subtopic $l$ being covered by the documents on the prior positions to $r_{i}$. Specifically, $c_{li}$ is defined as $c_{li}=\sum_{j:r_{j}\leq r_{i}}y_{jl}$.
Then, dividing the $\alpha\textrm{-}DCG$ of a given documents list by the $\alpha\textrm{-}DCG_{ideal}$ of an ideal ranking list yields the normalized $\alpha\textrm{-}NDCG$,
(2) $\alpha\textrm{-}NDCG=\frac{\alpha\textrm{-}DCG}{\alpha\textrm{-}DCG_{ideal}}$
The metric of top $k$ documents in a given list, called $\alpha-NDCG@k$, is usually used to measure the diversity of the documents list.
3.2.2. $ERR\textrm{-}IA$
The $ERR\textrm{-}IA$ is defined in Equation (3).
(3) $ERR\textrm{-}IA=\sum^{n}_{i=1}\frac{1}{r_{i}}\sum^{m}_{l=1}\frac{1}{m}\left(\prod_{j:r_{j}\leq r_{i}}(1-\frac{2^{y_{jl}}-1}{2^{y_{l}^{max}}})\right)\frac{2^{y_{jl}}-1}{2^{y_{l}^{max}}}$
If $y_{il}$ is a binary label ($y_{il}$ = 0 or 1), the Equation (3) can be rewritten as Equation (4).
(4) $ERR\textrm{-}IA=\sum^{n}_{i=1}\frac{1}{r_{i}}\sum^{m}_{l=1}\frac{1}{m}\frac{y_{il}}{2^{c_{li}+1}}$
Similar to $\alpha\textrm{-}NDCG@k$, $ERR\textrm{-}IA@k$ is also often used in practice.
3.2.3. $S\textrm{-}recall$
The $S\textrm{-}recall$ is always defined on top $k$ documents of a ranked list, i.e.,
(5) $S\textrm{-}recall@k=\frac{\left|\cup^{k}_{i=1}subtopics(d_{i})\right|}{\left|\cup^{n}_{i=1}subtopics(d_{i})\right|}$
Obviously, $S\textrm{-}recall@k$ represents the proportion of subtopics number recalled in the first top $k$ documents to the number of subtopics contained in the whole document list.
3.3. Test Collection for Diversified Search
Suppose there is a given query $\mathbf{q}$ which is associated with a set of candidate documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}\in\mathcal{D}$, where query $\mathbf{q}$ and each document $\mathbf{d_{i}}$ are represented as $L$-dimensional preliminary representations, i.e., the $L$-dimensional vector given by the BERT model (Devlin et al., 2018), and $\mathcal{D}$ is the set of all candidate documents. Our objective is to sort the documents in the candidate set $\mathcal{D}$ in such a way that the documents ranked higher cover as many subtopics as possible, thereby improving the diversity evaluation metrics.
To compute the diversity metrics introduced above, we need to collect a test collection with relevance labels at the subtopic level. Therefore, a test collection with $N$ labeled queries can be formulated as:
$\{(q^{k},D^{k}),J^{k}\}_{k=1}^{N}$
where $J^{k}$ is a binary matrix with $n\times m$ dimensions. $J^{k}(i,l)=1$ means that document $\mathbf{d_{i}}$ covers the $l$-th subtopic in the given query $q^{k}$ and $J^{k}(i,l)=0$ otherwise.
4. The MA4DIV Model
In this section, we first describe how to formulate the search results diversification problem as a Dec-POMDP by introducing the basic elements in the MA4DIV framework. Then, we present the architecture of the MA4DIV framework, as shown in Figure 1, and elaborate the implementation of the Agent Network, Ranking Process, and Mixing Network. Finally, we will give an algorithm to train the MA4DIV.
4.1. Essential Elements of MA4DIV
In Section 3.1, we define the process of diversity ranking as a Dec-POMDP. Next, we introduce the essential elements of Dec-POMDP specified to our MA4DIV:
Agents $\mathcal{G}$: We model each document $i\in\{1,\dots,n\}$ as agent $i\in\mathcal{G}\equiv\{1,\dots,n\}$. In the setting of multi-agent cooperation, each document implements the ranking process by cooperating with each other.
Observation $\mathcal{O}$: Observation represents the localized information that each agent receives. Each agent makes decisions based on its respective observation $\bm{o_{i}}$. The defination of $\bm{o_{i}}$ is shown as Equation (6).
(6) $\bm{o_{i}}=\left\{\mathbf{q},\mathbf{d_{i}},\textbf{D}\right\}$
where $\mathbf{q}$ is embedding vector of the given query, $\mathbf{d_{i}}$ is embedding vector of document $i$ and $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$ represents the embedding vectors of all documents.
State $\mathcal{S}$: State signifies global information and holds a pivotal role in multi-agent cooperation. It is through the utilization of state $\mathcal{S}$ that effective collaboration between agents (documents) can be established, resulting in a diverse and high-quality ranking list. State is defined $\bm{s}$ as:
(7) $\bm{s}$ =$=\{\mathbf{q},\textbf{D}\}$
From Equation (7) we can see that state $\bm{s}$ contains information for the given query $\mathbf{q}$ and information for all documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$.
Actions $\mathcal{A}$: The action space $a_{i}\in\mathcal{A}$ of the MA4DIV algorithm is defined as a set of discrete ranking scores $s_{i}\in\{1,\dots,\mathcal{|A|}\}$. When making decisions, agent $i$ will select an action $a_{i}$ from $\mathcal{A}$. The corresponding score $s_{i}$ serves as the basis for agent $i$ (document $i$) to assess the level of diversity to itself.
Upon the selection of joint-action $\mathbf{a}=\{a_{1},\dots,a_{n}\}$ by all agents based on joint-observation $\mathbf{o}=\{\bm{o_{1}},\dots,\bm{o_{n}}\}$, the ranking scores of candidate documents list $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$ can be obtained as $\mathbf{scores}=\{s_{1},\dots,s_{n}\}$. By sorting all documents according to $\mathbf{scores}$, a ranked document list by diversity level can be achieved, called $\mathbf{D_{ranked}}$.
Reward $\mathcal{R}$: The reward function should guide the parameter updates of the model to achieve diversity of ranked documents. In search result diversification, there are some evaluation metrics, such as $\alpha$-NDCG, are used to evaluate the diversity of a ranked document list. So it’s natural to consider these metrics as reward functions. In this paper, we define the reward function as $\alpha$-NDCG at top $k$ positions:
(8) $\mathcal{R}\left(\mathbf{D_{ranked}}\right)=\alpha\textrm{-}NDCG@k$
Episode $\mathcal{E}$: In typical MDP, an episode contains multiple decision steps. So the objective function is cumulative rewards in an episode with a discount factor $\gamma$, defined as:
(9) $G_{t}=\sum_{k=0}^{n-1-t}\gamma^{k}\mathcal{R}_{t+k+1}$
where $n$ is the step number in an episode.
However, the episode in MA4DIV contains only one step. Since the complete ranked document list can be obtained in one time step, the reward can then be calculated for end-to-end training. Therefore, instead of implementing a multi-step decision through a state transition function, an episode can contain only one step. And the final objective function of MA4DIV can be written as:
(10) $G=\mathcal{R}=\alpha\textrm{-}NDCG@k$
4.2. Agent Network
Figure 1. The Framework of MA4DIV.
In Figure 1, the green modules represent the Agent Network, which share parameters with each other. On the left side of Figure 1 is the detailed structure of the Agent Network. It can be seen that we concatenate $\{\mathbf{q},\mathbf{d_{i}},\mathbf{e_{i}}\}$, which is precisely the observation for agent $i$ defined in Equation (6). And $\mathbf{e_{i}}$ is a vector with high-level cross feature between document $i$ and all other documents, computed by Multi-Head Self-Attention (MHSA) module:
The MHSA module is an essential component in Transformer (Vaswani et al., 2017) and its variants. Given an input matrix $\mathbf{X}\in\mathbb{R}^{n\times L}$, where $n$ and $L$ denote the number of documents in candidiate document set D and the embedding vector dimension of $\mathbf{d_{i}}$. MHSA computes a weighted sum of all the input tokens for each token. The weight (also called attention score) between two tokens is calculated using their similarity.
Specifically, MHSA first linearly projects the input matrix into query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$ matrices, i.e., $\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{X}\mathbf{W}_{V}$, where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are learnable weight matrices with dimensions $L\times d_{k}$, $L\times d_{k}$, and $L\times d_{v}$ respectively, and $d_{k}$ and $d_{v}$ are the dimensions of keys/values. Then, the attention score between each two tokens is computed as $\mathbf{A_{s}}=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})$. The output of MHSA is $\mathbf{O}=\mathbf{A_{s}}\mathbf{V}$, with $\mathbf{O}\in\mathbb{R}^{n\times d_{v}}$.
MHSA further enhances the model capacity by applying the above process multiple times in parallel, which is called multi-head mechanism. Suppose there are $h$ heads, the output after multi-head is $\mathbf{O}=[\mathbf{O}_{1};\mathbf{O}_{2};\cdots;\mathbf{O}_{h}]\mathbf{W}_{O}$, where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation, and $\mathbf{O}_{i}$ and $\mathbf{W}_{O}$ are the output of the $i$-th head and a learnable weight matrix with dimensions $h\cdot d_{v}\times d$, respectively. Finally, $\mathbf{O}$ is reshaped to output $\mathbf{e_{i}}$ as Equation (11):
(11) $\{\mathbf{e_{1}},\dots,\mathbf{e_{i}},\dots,\mathbf{e_{n}}\}=\mathbf{MHSA}([\mathbf{d_{1}},\dots,\mathbf{d_{n}}])$
Why do we employ the MHSA module to obtain $\mathbf{e_{i}}$ rather than using other networks like MLP to compute the cross features of all documents directly? The main reason is as follows:
According to (Pang et al., 2020), the ranking model should satisfy the permutation invariance property which is described as following Defination:
Definition 0 (Permutation Invariance).
The target ranking for a given set is the same regardless of the order of documents in the set. In other words, no matter how exchange the positions of $d_{i}$ and $d_{j}$ in inputs information, it will not affect the final ranking result.
A method to construct a ranking model in accordance with Permutation Invariance involves initially attributing scores to the documents via a permutation equivariant scoring function, followed by ordering based on these scores. Pang et al. (2020) also proved that the multi-head attention block is permutation equivariant. Therefore, the MHSA module is used to compute higher-level cross feature vectors between documents.
Then, the action-values $Q_{i}^{a_{m}},m\in\{1,\dots,|\mathcal{A}|\}$ are calculated according to Equation (12). Specifically, $Q_{i}^{a_{m}}=Q(\bm{o_{i}},a_{m})$ represents the expected reward of taking a specific action $a_{m}$ in a given observation $\bm{o_{i}}$ for agent $i$, which is defined in Equation (13).
(12) $\{Q_{i}^{a_{1}},\dots,Q_{i}^{a_{m}},\dots,Q_{i}^{a_{n}}\}$ $=\mathbf{MLP}\left(\{\mathbf{q},\mathbf{d_{i}},\mathbf{e_{i}}\}\right)$
(13) $Q_{i}^{a_{m}}$ $=Q(\bm{o_{i}},a_{m})$
$=\mathbb{E}[\mathcal{R}|\bm{o}=\bm{o_{i}},a=a_{m}]$
In reinforcement learning (RL), a popular approach to select the action-value is the $\varepsilon$-greedy policy, which is used to balance exploration and exploitation during the learning process.
Specifically, with a probability of $1-\varepsilon$, the agent $i$ chooses the action-value $Q_{i}^{a_{*}}$ which has the highest estimated value ($\max Q_{i}^{a_{m}}$) for the current observation $\bm{o_{i}}$, and the corresponding chosen action is $a_{*}^{i}=\arg\max_{a_{m}}Q_{i}^{a_{m}},\forall m\in\{1,\dots,|\mathcal{A}|\}$. This is the exploitation part, where the agent makes the best decision based on the current model. On the other hand, with a probability of $\varepsilon$, the agent $i$ chooses an action-value uniformly at random from $\mathcal{|A|}$ action-values. This is the exploration part, where the agents try to discover new permutations of the candidate documents set. Mathematically, this process can be written as Equation (14):
(14) $Q_{i}^{a_{*}}=\begin{cases}\max Q_{i}^{a_{m}},\forall m\in\{1,\dots,|\mathcal{A}|\}&\text{with }p\textrm{=}1\textrm{-}\varepsilon\\ \text{random }Q_{i}^{a_{m}},\forall m\in\{1,\dots,|\mathcal{A}|\}&\text{with }p\textrm{=}\varepsilon\end{cases}$
In the training process, the value of $\varepsilon$ is often initially set to a high value (e.g., 1.0), and gradually decayed to a low value (e.g., 0.05) over training steps $t$ ($\varepsilon(t)=\max(0.05,1-\frac{t}{T})$), allowing the agent to explore the environment extensively at the beginning, and then exploit more as its knowledge increases. In the testing process, we set $\varepsilon=0$ to get the optimal policy.
In this way, the $\varepsilon$-greedy policy used in MA4DIV helps to balance the trade-off between exploration and exploitation, which promotes to explore more different ranking permutation. And it’s more conducive to update the parameters of the Agent Network.
4.3. Ranking Process
In the ranking process, the agents receive a query q and the associated documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$. As shown in Figure 1, each agent $i$ select its own action-value $Q_{i}^{a_{*}}$ and the corresponding action $a_{*}^{i}$ according to Equation (14). Because the action space $\mathcal{A}$ is defined as discrete ranking scores, that is, each action $a_{m}$ corresponds to a ranking score. So the ranking scores of all documents can be obtained as $\mathbf{scores}=\{s_{1},\dots,s_{n}\}$. Finally, a ranked document list $\mathbf{D_{ranked}}$ is obtained based on the descending order of $\mathbf{scores}$.
We show the pseudo-code of ranking process in Algorithm 1.
1Inputs: Give the query q and documents $\textbf{D}=\{\mathbf{d_{1}},\dots,\mathbf{d_{n}}\}$.
2 Select Actions: Each agent $i$ select action-value $Q_{i}^{a_{*}}$ to get $\{Q_{1}^{a_{*}},\dots,Q_{i}^{a_{*}},\dots,Q_{n}^{a_{*}}\}$ and $\{a_{*}^{1},\dots,a_{*}^{i},\dots,a_{*}^{n}\}$ base on Equation (14).
3 Getting Scores: Get the corresponding ranking scores $\mathbf{scores}=\{s_{1},\dots,s_{i},\dots,s_{n}\}$.
4 Sorting: Sort the candidate documents in D based on $\mathbf{scores}$.
Output: Get a ranked document list $\mathbf{D_{ranked}}$.
Algorithm 1 The Ranking Process of MA4DIV
4.4. Value-Decomposition for MA4DIV
After the ranking process, the reward $\mathcal{R}$ is obtained according to Equation (8) based on $\mathbf{D_{ranked}}$. Obviously, $\mathcal{R}$ is a scalar while there are $n$ action-values $\{Q_{1}^{a_{*}},\dots,Q_{n}^{a_{*}}\}$ for all agents. $\mathcal{R}$ is the evaluation of the global ranking effect, while $Q_{i}^{a_{*}}$ is related to the score of a single document. Therefore, we cannot use $\mathcal{R}$ to directly update the action-value functions of all agents.
The MARL algorithms based on value decomposition, such as (Sunehag et al., 2017; Rashid et al., 2018; Wang et al., 2020b), consider that there is the global utility function $Q_{tot}^{*}$ can be decomposed into the action-value functions $Q_{i}^{a_{*}}$ of all individual agents. $Q_{tot}^{*}$ represents the action-value utility of the joint-action $\mathbf{a}$ under the joint-observation information $\mathbf{o}$, and is also a scalar value. The monotonic decomposition relation can be expressed as Equation (15):
(15) $\arg\max_{\mathbf{a}}Q_{tot}^{*}(\mathbf{o},\mathbf{a})=\begin{pmatrix}\arg\max_{{a}_{1}}Q_{1}(\bm{o_{1}},{a}_{1})\\ \vdots\\ \arg\max_{{a}_{n}}Q_{n}(\bm{o_{n}},{a}_{n})\end{pmatrix}$
In online ranking process, all agents adopt a greedy strategy to score documents, which is consistent with the maximization operation on the right side of Equation (15). So we just need to learn how to maximize $Q_{tot}^{*}$. Since $Q_{tot}^{*}$ is a global utility function and $\mathcal{R}$ is also used to measure the global ranking effect, we can maximize $Q_{tot}^{*}$ with $\mathcal{R}$, thus maximizing the utility of each individual agent ($Q_{i}^{a_{*}}$). The specific loss function will be introduced in the following Section 4.6 (Equation (19)).
Furthermore, QMIX (Rashid et al., 2018) proposes a sufficient but not necessary realization to satisfy Equation (15). Monotonicity can be enforced by a constraint condition on the relationship of partial derivative between $Q_{tot}^{*}$ and $Q_{i}^{a_{*}}$:
(16) $\frac{\partial Q_{tot}^{*}}{\partial Q_{i}^{a_{*}}}\geq 0,\hskip 8.5359pt\forall i\in\mathcal{G}\equiv\{1,\dots,n\}$
Additionally, in Equation (15) and (16), it is established that $Q_{i}^{a_{*}}=\max Q_{i}(\bm{o_{i}},{a}_{m}),a_{*}^{i}=\arg\max_{a_{m}}Q_{i}(\bm{o_{i}},{a}_{m}),\forall m\in\{1,\dots,|\mathcal{A}|\}$.
4.5. Mixing & Hyper Network Structure
To convert the individual action-value functions $Q_{i}^{a_{*}}$ into the global utility function $Q_{tot}^{*}$ and ensure compliance with Equation (16), an architecture consisting of a mixing network and a set of hypernetworks is designed. The right part of Figure 1 illustrates the detail structure.
Mixing network, a fully connected network, utilizes the action-values $\{Q_{1}^{a_{*}},\dots,Q_{n}^{a_{*}}\}$ from the agent network as inputs, and then, monotonically combines these inputs to generate $Q_{tot}^{*}$. The calculation process is shown as:
[…]
Hypernetworks, several fully connected networks, play a role in generating mixing network weights. Firstly, each hypernetwork takes global state $\bm{s}$ as input and generate the vectors of $\mathbf{W1}$, $\mathbf{B1}$, $\mathbf{W2}$, B2 respectively. Then, the vectors are reshaped into matrices of appropriate size as parameters for the mixing network. Lastly, to achieve the monotonicity constraint of Equation (16), $\mathbf{W1}$, $\mathbf{W2}$ are limited to non-negative by an absolute activation function. However, biases $\mathbf{B1}$, B2 do not need to be set to non-negative because biases has nothing to do with the partial derivatives between $Q_{tot}^{*}$ and $Q_{i}^{a_{*}}$. The hypernetworks operate as follows:
(18) $\mathbf{W1}&=\mathbf{Reshape}\left(\left|\mathbf{FC_{{\scriptscriptstyle W1}}}\left(\bm{s}\right)\right|\right),\mathbf{B1}=\mathbf{Reshape}\left(\mathbf{FC_{{\scriptscriptstyle B1}}}\left(\bm{s}\right)\right)\\ \mathbf{W2}&=\mathbf{Reshape}\left(\left|\mathbf{FC_{{\scriptscriptstyle W2}}}\left(\bm{s}\right)\right|\right),\text{B2}=\mathbf{Reshape}\left(\mathbf{FC_{{\scriptscriptstyle B2}}}\left(\bm{s}\right)\right)\end{aligned}\end{gathered}$
4.6. Training Process of MA4DIV
In this section, we introduce the complete training process for MA4DIV. The parameters of models are $\Theta=\{\theta,\psi,\phi\}$ where $\theta$ represents the parameters of agent network, $\psi$ and $\phi$ are the parameters of mixing network and hypernetworks respectively. The model parameters $\Theta$ is updated on $N$ training data with the form of $\{(q^{k},D^{k}),J^{k}\}_{k=1}^{N}$.
Algorithm 2 shows the training pseudo-code. At the beginning, parameters $\Theta$ are initialized randomly and the replay buffer $\mathcal{M}$ is set to $\varnothing$. A whole training epoch consists of two parts. The first part is the stage of data generation. Given data $\mathbf{q}=q^{i}$, $\mathbf{D}=D^{i}$, with subtopic labels $J^{i}$, tuple $\mathcal{T}$ in the episode is obtained based on the definition in Equation (6, 7, 8). Then, store the tuple $\mathcal{T}$ in replay buffer. Each epoch can generate a kind of permutation of all documents $\{D^{k}\}_{k=1}^{N}$ in all training data. The second part is updating models with parameters $\Theta$ for $N\_update$ times. And the update is performed by sampling data through a mini-batch approach from $\mathcal{M}$. The whole network is trained end-to-end by minimizing the loss shown in Equation (19) and (20) to maximize $Q_{tot}$.
(19) $\mathcal{L}_{TD}=\sum_{i=1}^{b}\left(y_{i}^{tot}-Q_{tot}(\textbf{o},\textbf{a},\bm{s};\Theta)\right)^{2}$
(20) $y^{tot}$ $=\mathcal{R}+\gamma\max_{a^{\prime}}Q_{tot}(\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime};\Theta^{\prime})$
$=\mathcal{R}$
1 Initialize: The parameters of network $\Theta=\{\theta,\psi,\phi\}$, replay buffer $\mathcal{M}$.
2 Inputs: Training data $\{(q^{k},D^{k}),J^{k}\}_{k=1}^{N}$.
3
4for $epoch=1$ to $N\_epoch$ do
5       // generate data
6       for $i=1$ to $N$ do
7             Current training data: $\mathbf{q}=q^{i}$, $\mathbf{D}=D^{i}$, with subtopic labels $J^{i}$.
8             Rollout an episode following Algorithm 1.
9             Collect a tuple $\mathcal{T}$ = ($\mathbf{o},\bm{s},\mathbf{a},\mathcal{R},\mathbf{o^{\prime}},\bm{s^{\prime}}$) during this episode.
10             Store the tuple $\mathcal{T}$ in $\mathcal{M}$.
11
12
13      // update models
14       for $update=1$ to $N\_update$ do
15             Sample a random minibatch $\bm{b}$ from $\mathcal{M}$.
16             Calculate $y^{tot}$ and loss $\mathcal{L}$ for all sampled data from $\bm{b}$ based on Equation (20) and (19).
17             Update the parameters of networks $\Theta$ by gradient descent.
18
19
20
21Ouput: A well-trained ranking model with parameters $\Theta$.
22
Algorithm 2 The Training Process of MA4DIV
Table 1. Comparisons of datasets TREC and DU-DIV.
Dataset Source No. of Queries No. of Documents per Query No. of Subtopics Representation
TREC TREC 2009$\sim$2012 Web Track 198 average 211 docs/query maximum 7 subtopics/doc Doc2vec (Le and Mikolov, 2014)
DU-DIV Baidu 4473 15 docs/query total 50 subtopics BERT (Devlin et al., 2018)
In Equation (19), $b$ is the batch size of sampled data from replay buffer. In Equation (20), $Q_{tot}(\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime};\Theta^{\prime})$ is calculated based on information of the next time step, i.e., $\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime}$, in typical MDP which contains multiple decision steps. However, we define that MA4DIV only have one time step for an episode, so $\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime}$ do not exist and $Q_{tot}(\textbf{o}^{\prime},\textbf{a}^{\prime},\bm{s}^{\prime};\Theta^{\prime})$ equals to 0, that is $y^{tot}=\mathcal{R}$.
Table 2. Hyperparemeters of MA4DIV on different dataset.
Dataset $|\mathcal{A}|$ $L$ $H$ $z$
TREC 30 1 4 64
DU-DIV 15 1 4 256
5. Experiments
Our experiments mainly focus on next research questions:
(RQ.1) How does MA4DIV perform on the TREC Web Track datasets when compared to existing diversified search baselines?
(RQ.2) How does MA4DIV perform on a large-scale industrial dataset comparing to existing baselines?
(RQ.3) How efficient is MA4DIV in training process?
5.1. Datasets
We conduct experiments to address above research questions on two datasets, TREC 2009$\sim$2012 Web Track datasets and DU-DIV dataset. TREC 2009$\sim$2012 Web Track datasets are publicly available and Xia et al. (2017) first uses them for training and evaluating diversified search models. Almost all the following works, such as $\mathrm{M^{2}Div}$ (Feng et al., 2018), DALETOR (Yan et al., 2021), MO4SRD (Yu, 2022), etc., conduct experiments on these datasets. As a widely used test collection for diversified search, the TREC Web track dataset only contains 198 queries, and previous studies often use Doc2vec (Le and Mikolov, 2014) to obtain the document embeddings on this dataset. In this work, to ensure a fair and meaningful comparison with existing work, we first compare MA4DIV with existing baselines on the this relatively small dataset. Then, we further conduct extensive experiments on a new and larger scale diversity dataset called DU-DIV, which is constructed based on real user data in Baidu search engine.
The details of these two datasets are shown in Table 1. The DU-DIV dataset has 4473 queries, far more than the TREC datasets. The documents in the DU-DIV dataset are the top-15 documents retrieved by the Baidu search engine for each query, which makes this dataset particularly useful for the diversified ranking task in which we need to rerank these top documents for diversity. In the TREC web track datasets, each document contains a maximum of 7 subtopics. In the DU-DIV dataset, we predefined 50 subtopics, such as law, healthcare, education, etc. Each document corresponds to one or more of the predefined subtopics. Moreover, while previous studies on the TREC web track dataset often use the Doc2vec model to represent queries and documents in vector with $L\textrm{=}100$ dimensions, the DU-DIV dataset utilizes the more powerful model BERT (Devlin et al., 2018) to encode queries and documents into vectors with $L\textrm{=}1024$ dimensions.
5.2. Experimental Settings
Considering that TREC datasets contain only 198 queries, this is prone to overfitting for existing deep learning models. In addition, the average number of documents related to each query is about 211, but a large number of these documents do not contain any subtopics. Therefore, we reshape TREC dataset. That is, from each of the original document lists in TREC (198 lists in total, with an average of 211 documents per query), we subsample 30 documents at a time to form a new document list. This process runs multiple times over a document list associated with the same query, resulting in a reshaped dataset with a total of 6,232 document lists (198 unique queries). And it’s crucial to ensure that an appropriate number of the 30 documents sampled contain at least one subtopic.
We conduct 5-fold cross-validation experiments on the reshaped TREC dataset with the same subset split based on unique queries as in (Feng et al., 2018). In the DU-DIV dataset, we randomly sample 80% of 4473 unique queries and their associated document lists to the training set, and the left 20% to the test set.
The code framework of MA4DIV is based on PyMARL2 framework (Hu et al., 2021), which is a open-source code framework for multi-agent reinforcement learning. The key hyperparameters of MA4DIV on two datasets are shown in Table 2. $|\mathcal{A}|$ is the dimension of action space, i.e. the number of discrete scores (1, …, $\mathcal{|A|}$). $L$, $H$ and z respectively represent the number of MHSA block, attention heads and attention embedding dimensions of MHSA.
Table 3. Performance comparison of baselines and MA4DIV on TREC web track datasets. The best result in each metric is bold and the top-3 results are shaded. ”*” indicates the difference between the baseline and M4DIV (ours) is statistically significant.
Method $\alpha$-NDCG@5 $\alpha$-NDCG@10 ERR-IA@5 ERR-IA@10 S-recall@5 S-recall@10
MMR 0.4273* 0.5059* 0.2015* 0.2153* 0.6058* 0.7848*
xQuAD 0.4451* 0.5296* 0.2108* 0.2243* 0.5994* 0.7887*
MDP-DIV 0.4987* 0.5663* 0.2662 0.2865 0.6310* 0.7853*
$\mathrm{M^{2}Div}$ 0.5144* 0.5798* 0.2338* 0.2448* 0.6593 0.8028
DALETOR 0.5008 0.5703 0.2237* 0.2355* 0.6451* 0.8012
MO4SRD 0.5135* 0.5815* 0.2280* 0.2396* 0.6576 0.8074
MA4DIV (ours) 0.5056 0.5724 0.2680 0.2882 0.6600 0.8070
We compare MA4DIV with several state-of-the-art baselines in search result diversification, including the heuristic methods:
MMR (Carbonell and Goldstein, 1998): a heuristic approach which select document according to maximal marginal relevance.
xQuAD (Santos et al., 2010): a representative approach that explicitly models different aspects underlying the query in the form of subqueries.
MDP-DIV (Xia et al., 2017): a single-agent reinforcement learning approach which model the diverse ranking process as MDP.
$\mathbf{M^{2}}$iv (Feng et al., 2018): a single-agent reinforcement learning approach which utilizes Monte Carlo Tree Search to enhance ranking policy.
DALETOR (Yan et al., 2021): a method for approximating evaluation metrics that optimizes an approximate and differentiable objective function.
MO4SRD (Yu, 2022): another method to make evaluation metrics differentiable and optimize the approximating evaluation metrics.
We reproduce the code of DALETOR and MO4SRD, and ensure that these two algorithms have the same parameter scale of MHSA structure as our MA4DIV in Table 2, that is, the $L=1$ and $H=4$, $z$ is 64 and 256 respectively on two datasets.
5.3. Experimental Results
Table 3 and 4 report the performances of all baselines and our MA4DIV in terms of the six diversity performance metrics, including $\alpha$-NDCG@5, $\alpha$-NDCG@10, ERR-IA@5, ERR-IA@10, S-recall@5, S-recall@10 on the TREC web track datasets and the DU-DIV dataset respectively. We highlight the top three algorithms on each metric by darkening the background color of the numbers.
5.3.1. TREC web track datasets
In Table 3, $\mathrm{M^{2}Div}$ and MO4SRD have the best performance in term of $\alpha$-NDCG, our MA4DIV performs worse than these two algorithms but better than the others. Moreover, MA4DIV takes the lead in both ERR-IA and S-recall metrics, with MO4SRD also exhibiting strong performance in S-recall metric. Considering multiple evaluation metrics comprehensively, MA4DIV, $\mathrm{M^{2}Div}$ and MO4SRD are the best three algorithms on the TREC web track datasets. (answer to RQ.1)
Table 4. Performance comparison of baselines and MA4DIV on DU-DIV dataset. The best result in each metric is bold and the top-3 results are shaded. ”*” indicates the difference between the baseline and M4DIV (ours) is statistically significant.
Method $\alpha$-NDCG@5 $\alpha$-NDCG@10 ERR-IA@5 ERR-IA@10 S-recall@5 S-recall@10
MMR 0.7416* 0.8171* 0.7583* 0.8100* 0.6062* 0.8481*
xQuAD 0.7075* 0.8277* 0.7745* 0.8201* 0.5670* 0.8520*
MDP-DIV 0.8101* 0.8623 0.7694* 0.8204* 0.6389* 0.8590*
$\mathrm{M^{2}Div}$ 0.8089* 0.8610* 0.7633* 0.8142* 0.6468 0.8608*
DALETOR 0.7981* 0.8543* 0.7450* 0.7955* 0.6447* 0.8550*
MO4SRD 0.7989* 0.8553* 0.7467* 0.7969* 0.6432* 0.8556*
MA4DIV (ours) 0.8187 0.8702 0.7836 0.8342 0.6534 0.8699
As mentioned in Section 5.1, TREC web track datasets have only 198 unique queries, deep learning methods are prone to overfitting on this dataset. Figure 2 shows the training curve of MA4DIV on the TREC web track datasets. We can see, after about the 11 iterations, MA4DIV achieves the best performance on the test data, and the subsequent iterations fall into overfitting the relatively small training set. However, the evaluation metric $\alpha$-NDCG@10 on the training set grows to more than 0.8. These results indicate that:
  • Our MA4DIV is capable to effectively optimize the diversity metric $\alpha$-NDCG@10.
  • However, it could easily overfit the relatively small training set of the TREC web track dataset, which only contains 198 unique queries.
  • Therefore, MA4DIV has the potential to perform even better in the search result diversification task when given a larger dataset.
Therefore, we further construct the DU-DIV dataset and conduct extensive experiments on it.
Figure 2. Training curves of MA4DIV on TREC web track datasets.
5.3.2. DU-DIV dataset
Table 4 shows results in different evaluation metrics on the DU-DIV dataset. Obviously, MA4DIV achieves state-of-the-art performances on all six evaluation metrics. Heuristic algorithms, MMR and xQuAD, perform poorly on $\alpha$-NDCG. Comparing to single-agent reinforcement learning methods, MDP-DIV and $\mathrm{M^{2}Div}$, MA4DIV has a better performance on evaluation metrics which demonstrate that MA4DIV can improve the diversity of search results. And MA4DIV also outperforms DALETOR and MO4SRD which optimize the approximate value of the evaluation metrics. This verifies that directly optimizing the evaluation metrics is indeed better than optimizing the approximation of the evaluation metrics. All these results demonstrate that the proposed MA4DIV method outperforms existing search result diversification methods on an industrial scale dataset (answering RQ.2).
Interestingly, we notice that MDP-DIV, which does not perform well on the TREC web track datasets, is the second best performing algorithm on the DU-DIV dataset, following our MA4DIV in terms of $\alpha$-NDCG@5 and $\alpha$-NDCG@10. In MDP-DIV, the reward function is defined as: $R(s_{t},a_{t})=\alpha\textrm{-}DCG[t+1]-\alpha\textrm{-}DCG[t]$, where $\alpha\textrm{-}DCG[t]$ is the discounted cumulative gain at the $t$-th ranking position. And the objective function of MDP-DIV is to maximize the cumulative rewards in one episode, that is $G_{t}=\sum_{k=0}^{n-1-t}\gamma^{k}R_{t+k+1}$. Consider that Xia et al. (2017) sets $\gamma\textrm{=}1$ and $\alpha\textrm{-}DCG[t]\textrm{=}0$, the objective function can be rewritten as $G_{t}=\alpha\textrm{-}DCG[n]$, which is equal to $\alpha\textrm{-}DCG$ defined in Equation (2). In addition, the objective function of our MA4DIV defined in Equation (10) is equal to $\alpha\textrm{-}NDCG$. Essentially, the optimization objective of MDP-DIV is basically the same as that of our MA4DIV. Therefore, due to the correctness of its optimization objective, MDP-DIV can outperform all the other baselines.
5.4. Comparison and Analysis of Efficiency
Table 5. The comparison of the training time taken by different algorithms to achieve the optimal value of metrics.
Method Training Time
MDP-DIV 5$\sim$6 hours
MO4SRD 6$\sim$7 hours
$\mathrm{M^{2}Div}$ 1 day +
MA4DIV (ours) 20$\sim$25 mins
5.4.1. Training Time on TREC web track datasets
To highlight the efficiency of our MA4DIV, we also compare several algorithms from the perspective of training time. Table 5 shows the comparison of training time taken by different algorithms to achieve their optimal performance on the TREC web track datasets. To ensure a fair comparison, all the algorithms in Table 5 run on the same machine. We can see that $\mathrm{M^{2}Div}$ takes over a day to reach optimal performance. This is because $\mathrm{M^{2}Div}$ utilizes Monte Carlo Tree Search for policy enhancement, which is highly time-consuming. The time taken by MDP-DIV and MO4SRD is close and significantly less than $\mathrm{M^{2}Div}$. The proposed MA4DIV takes the shortest time, approximately only 20$\sim$25 minutes. This is due to the efficient exploration ability of using a multi-agent to model the ranking process, in which a permutation of the given documents set ($n$ documents) can be obtained at one time step. The single agent reinforcement learning algorithm MDP-DIV can only explore a permutation of the given documents after a whole episode which contains $n$ time steps. This is the main reason why the training time of MA4DIV is significantly shorter than MDP-DIV. In MO4SRD, the complexity of calculating all approximate ranking positions $r_{i}$ in Equation (1) is $\mathcal{O}(n^{2})$ and the complexity of calculating all $c_{li}$ approximations is $\mathcal{O}(n^{2}m)$, which are time-consuming parts. While MA4DIV directly optimizes Equation (2) without complex approximations, which takes significantly less time.
In summary, the training time of MA4DIV is significantly shorter than that of other baselines, which demonstrate its training efficiency and answers RQ.3.
5.4.2. Iteration Times on DU-DIV dataset
As the optimization goals of MA4DIV and MDP-DIV are essentially the same as we discussed in Section 5.3.2, we are interested in comparing the learning efficiency of the proposed multi-agent RL approach and the existing single-agent RL method.
Figure 3 shows the training curves of MA4DIV and MDP-DIV on the DU-DIV dataset. We conduct the experiments of these two algorithms with the same learning rate which is set to $1\times 10^{-5}$. From the curves, we can see that, after 250 iterations, $\alpha$-NDCG of MA4DIV basically converges to about 0.87. However, $\alpha$-NDCG of MDP-DIV still fails to reach the optimal value at around 6000 iterations. This indicates that MA4DIV has higher exploration and exploitation efficiency compared with MDP-DIV, which also answers RQ.3.
The main reason for the results just mentioned is as follows: MDP-DIV adopts an on-policy RL method to optimize the cumulative rewards. A feature of on-policy RL is that the currently sampled data is discarded after it is utilized. While MA4DIV is a off-policy RL method. From 9-th line and 12-th line of Algorithm 2, we can see that the sampled data is stored in replay buffer $\mathcal{M}$, then update the model multiple times in minibatch approach. In addition, each data stored in replay buffer $\mathcal{M}$ may be sampled multiple times to update the ranking model. Therefore, the data utilization efficiency of MA4DIV is higher than that of MDP-DIV, which is the main reason why MA4DIV converges significantly faster.
The shorter training time on the TREC web track datasets and the faster convergence rate on the DU-DIV dataset jointly demonstrate that MA4DIV has high efficiency in training, which answers RQ.3.
Figure 3. Training curves of MA4DIV and MDP-DIV on DU-DIV dataset.
6. Conclusions and Future Work
In this paper we have proposed a new method for search result diversification, referred to as MA4DIV. To alleviate the sub-optimal problem and improve the efficiency of training, we consider each document as an agent and model the diversity ranking process as a multi-agent cooperative task. We conduct experiments on the TREC web track dataset to prove the effectiveness and potential of our MA4DIV. Furthermore, we construct a large-scale dataset called DU-DIV using the real data from Baidu search engine. And MA4DIV achieve state-of-the-art performance and high learning efficiency on the DU-DIV dataset.
A major contribution of this paper is to introduce a multi-agent reinforcement learning framework for the ranking optimization task. We argue that modeling ranking process in the multi-agent setting can be used not only to optimize the diversity of search results, but also to optimize other ranking tasks, such as relevance ranking and fairness ranking. Therefore, in future work, we will try to propose a general ranking framework base on multi-agent reinforcement learning and apply it to a variety of ranking tasks.